OpenStudy (anonymous):

what is the power output of an engine thaat does 60 000 j of work in 10 s?

OpenStudy (anonymous):

power=60000/10=6000W

OpenStudy (anonymous):

use the relation power=work/time

OpenStudy (anonymous):

power output of an engine

OpenStudy (anonymous):

same thing.

OpenStudy (anonymous):

yes. 6000Watts, 6kWatts or even 8.04613254

OpenStudy (anonymous):

8.04hp

OpenStudy (anonymous):

ic same thing from the last question im slow hahha

OpenStudy (anonymous):

\[Power = \ \frac{Work}{time} \ = \ \frac{1 Joule}{1 second} \ = 1 Watt\] Also the relationship between Watts and Horsepower(hp) is :\[1 horsepower = 745.699872 watts\]

OpenStudy (anonymous):

ic